OpenStudy (anonymous):

im doing cooridinate Geometry and needed help with Graphing Straight lines y=x+2 x|0|?| y| ?|0| i need working out please

OpenStudy (imstuck):

What is this: x|0|?| y| ?|0|

OpenStudy (anonymous):

it a graph

OpenStudy (imstuck):

What exactly are you wanting to do with all that up there?

OpenStudy (anonymous):

trying to find what is X and what is y

OpenStudy (imstuck):

ok hold on, I think I can help you...

OpenStudy (imstuck):

Do you want to make an x/y chart to graph that line?

OpenStudy (lochana):

x=-2 , y=2

OpenStudy (anonymous):

lochana i need the working out for it

OpenStudy (lochana):

well, if x=0 then y= 0 +2 which is 2 , if y =0 then 0 = x +2 which is x = -2

OpenStudy (anonymous):

thanks man